OpenStudy (anonymous):

Find the dimensions and volume of the largest cylinder that can be packed inside a box that has dimensions 14in. (width) by 7in. (width) by 2in. (lenght) Explain your reasoning. How much extra space would there be in the box?

OpenStudy (anonymous):

radium of the cylinder is 7/2=3.5in height: 2in pi x 3.5^2 x 2 = 76.97 This is the biggest cylinder you can fit inside

OpenStudy (anonymous):

extra space: 14x7x2-76.97=119.03